# Gaussian periods and Shanks’ cubic polynomials. II

### v1

@@ 3. Structure of the units group of the associated order
(3.2) $\mathcal{O}_{L}=\begin{cases}\mathcal{A}_{L/\mathbb{Q}}{\rm Tr}_{\mathbb{Q}(\mathfrak{f})/L}(\zeta_{\mathfrak{f}})&\text{if $p\nmid\mathfrak{f}$,}\\ \mathcal{A}_{L/\mathbb{Q}}({\rm Tr}_{\mathbb{Q}(\mathfrak{f})/L}(\zeta_{\mathfrak{f}})+1)&\text{if $p|\mathfrak{f}$}.\end{cases}$
@@
From this result, we know that if $\eta$ is a generator of $\mathcal{A}_{L/\mathbb{Q}}$-module $\mathcal{O}_{L}$, then there exists $u\in\mathcal{A}_{L/\mathbb{Q}}^{\times}$ such that $\eta=u{\rm Tr}_{\mathbb{Q}(\mathfrak{f})/L}(\zeta_{\mathfrak{f}})$ (resp. $\eta=u({\rm Tr}_{\mathbb{Q}(\mathfrak{f})/L}(\zeta_{\mathfrak{f}})+1)$) if $p\nmid\mathfrak{f}$ (resp. $p|\mathfrak{f}$), and there is a one-to-one correspondence between the set of all generators of $\mathcal{O}_{L}$ and $\mathcal{A}_{L/\mathbb{Q}}^{\times}$. In this section, we consider the structure of the group $\mathcal{A}_{L/\mathbb{Q}}^{\times}$.
@@ 5. Examples
We consider cyclic cubic fields with conductor $\mathfrak{f}=9\times 7\times 13$. All the pairs $(M,N)$ satisfying (1.5) are $(M,N)=(-3\cdot 19,1),(3\cdot 17,5),(3\cdot 8,10),(-3,11)$, and the corresponding pair $(n_{1},n_{2})$ are (30,1),(18,5),(3,10),(18,11)(-30,1),(18,5),(-3,10),(-18,11) in order. Table 1 shows the Gaussian periods of $L_{n}$, the period polynomials and Shanks’ cubic polynomials for each $n=n_{1}/n_{2}$.

### v3

@@ 3. Structure of the units group of the associated order
(3.2) $\mathcal{O}_{L}=\begin{cases}\mathcal{A}_{L/\mathbb{Q}}{\rm Tr}_{\mathbb{Q}(\zeta_{\mathfrak{f}})/L}(\zeta_{\mathfrak{f}})&\text{if $p\nmid\mathfrak{f}$,}\\ \mathcal{A}_{L/\mathbb{Q}}({\rm Tr}_{\mathbb{Q}(\zeta_{\mathfrak{f}})/L}(\zeta_{\mathfrak{f}})+1)&\text{if $p|\mathfrak{f}$}.\end{cases}$
@@
From this result, we know that if $\eta$ is a generator of $\mathcal{A}_{L/\mathbb{Q}}$-module $\mathcal{O}_{L}$, then there exists $u\in\mathcal{A}_{L/\mathbb{Q}}^{\times}$ such that $\eta=u{\rm Tr}_{\mathbb{Q}(\zeta_{\mathfrak{f}})/L}(\zeta_{\mathfrak{f}})$ (resp. $\eta=u({\rm Tr}_{\mathbb{Q}(\zeta_{\mathfrak{f}})/L}(\zeta_{\mathfrak{f}})+1)$) if $p\nmid\mathfrak{f}$ (resp. $p|\mathfrak{f}$), and there is a one-to-one correspondence between the set of all generators of $\mathcal{O}_{L}$ and $\mathcal{A}_{L/\mathbb{Q}}^{\times}$. In this section, we consider the structure of the group $\mathcal{A}_{L/\mathbb{Q}}^{\times}$.
@@ 5. Examples
We consider cyclic cubic fields with conductor $\mathfrak{f}=9\times 7\times 13$. All the pairs $(M,N)$ satisfying (1.5) are $(M,N)=(-3\cdot 19,1),(3\cdot 17,5),(3\cdot 8,10),(-3,11)$, and the corresponding pair $(n_{1},n_{2})$ are (30,1),(18,5),(3,10),(18,11)(-30,1),(18,5),(-3,10),(-18,11) in order. Table 1 shows the Gaussian periods of $L_{n}$, Shanks’ cubic polynomials, and the period polynomials for each $n=n_{1}/n_{2}$. Put $\rho=\rho_{n}$.